Sol Summary Report – February 22th

Crew 275 Sol Summary Report 22Feb2023

Sol: 10

Summary Title: Wuthering Heights and Rainy Nights

Author’s name: Jeremy Rabineau, Commander.

Mission Status: Nominal.

Sol Activity Summary:

Since yesterday evening, our activities have been highly impacted by the weather conditions. The noise of the wind was very impressive and impaired the sleep of some crewmembers. We were all very afraid for our atmospheric experiments, especially for the LOAC instrument that is very sensitive to rain. In the morning, we were authorized to leave the Hab for EVA #8 and stopped on our way to check the status of our instruments, as per EVA request. We noticed that both the meteorological mast and the LOAC had toppled over. We only retrieved the LOAC and left the meteorological mast, because we were aware that the wind speed would remain quite high. The crew of EVA #8 then reached North Ridge to perform their main mission, but after about 40 minutes they decided to come back because they were too cold, and the wind speed was too high to safely conduct their operations.

Following this event, we decided to also cancel EVA #9 that was planned for the afternoon.

On the campus, the environmental sensors have been fully deployed on all the buildings and can now give us information about temperature, humidity, pressure, and luminosity.

Look Ahead Plan:

We had to adapt our plans because of the weather. We will wait for better wind conditions to put back the antenna of Mega-Ares back at its initial place.

Today, two crewmembers have trained with the 3D map of North Ridge obtained by photogrammetry. Tomorrow, we would like to test their ability to reach some specific checkpoints there, as compared to the EVA crew of today (using only a 2D map). We are thus planning an EVA tomorrow morning (Sol 11), should the weather conditions allow it.

Anomalies in work: None.

Weather: Very cloudy and very windy,low 0°C/32°F, high 9°C/48°F.

Crew Physical Status: Optimal.

EVA: Because of the bad weather conditions, the team of EVA#8 had to retrieve the LOAC instrument, which was done successfully. They had to shorten their EVA because of high wind, which means that the scientific objectives were only partially achieved. EVA #9 was cancelled because the LOAC instrument was already retrieved, and the weather conditions were not good enough to safely leave the Hab.

Reports to be filed: Sol Summary, EVA Request, EVA Report, Operations Report, Journalist Report, GreenHab Report, HSO Report.

Support Requested: None.

Copyright © The Mars Society. All rights reserved. | Main Site